Find the pH and percent ionization of a 0.100 M solution of a weak monoprotic acid having the given Ka values.

(a) Ka = 1.9 10-5

(b) Ka = 1.9 10-3

(c) Ka = 1.9 10-1
Answer #1

We know that the equilibrium of a weak monoprotic acid is as follows:

HA <-> H+ + A-

We have 0.1 M HA, so using the constant definition:

Ka = [H+][A-] / [HA]

Making an ICE table:

HA <-> H+ + A-
I 0.1 M 0 M 0 M
C -x +x +x
E 0.1 - x x x

For letter a):

1.9 x 10-5 = x2 / (0.1 - x)

Isolating for x:

x = [H+] = 0.00136894

pH = -log (0.00136894) = 2.8636

% ionization = (0.00136894/0.1) * 100 = 1.3689%

For letter b):

1.9 x 10-3 = x2 / (0.1 - x)

Isolating for x:

x = [H+] = 0.0128667

pH = -log (0.0128667) = 1.89

% ionization = (0.0128667/0.1) * 100 = 12.8667%

For letter c)

1.9 x 10-1 = x2 / (0.1 - x)

Isolating for x:

x = [H+] = 0.0724

pH = -log (0.0724) = 1.14

% ionization = (0.0724/0.1) * 100 = 72.4%